Amore:

The sum of the present ages of two children is 18. Six years from now the age of the older child will the twice the age of the younger child. Find their present ages. Show the solution.

Mercury:

let's suppose y = the age of the younger child, and o = the age of the older child if their ages *now* sum up to 18, then y + o = 18 now, six years from now, their ages will be y + 6 and o + 6. since it says "six years from now... the older child will be twice the age of the younger child" ---> o + 6 = 2(y + 6)

Mercury:

so your system so far is: y + o = 18 o + 6 = 2(y + 6) you could solve this with either substitution or elimination. Iif you use substitution method: solve the first equation for y (alternatively you could solve for o, but I'll solve for y here) if y + o = 18 subtracting o from both sides: then y = 18 - o substitue "18 - o" for y in the second equation and solve for o. then go back to y + o = 18, plug in the o-value, and solve for y.

Amore:

Wow. Thank you ❤️❤️

Amore:

Can you help me with this problem too?? Three volumes of the series Mathematics: Its Content, Methods, and Meaning are on a shelf with no space between the volumes. Each volume is 1 inch thick without its covers. Each cover is 1/8 inch thick. A bookworm bores horizontally from the first page of Volume 1 to the last page of Volume III. How far does the bookworm travel?

Mercury:

each book has a front cover, the pages, and a back cover volume 1: the worm starts at page 1 (which is ahead of the front cover) so the worm only goes through the pages (1 inch) + the back cover (1/8) inch volume 2: the worm starts at the front cover and goes all the way to the end of the back cover. so it travels 1/8 inch + 1 + 1/8 inch volume 3: it stops at the last page, so right before the back cover. so it only travels through the front cover + pages. ----> 1/8 inch + 1 inch add up the three distances

Amore:

Thank yo so much.